OpenStudy (anonymous):

An arrow travelling at 45 m/s strikes and embeds itself in a 450 g apple which is initially at rest. They move off horizont ally at 12 m/s after impact. What is the mass of the arrow?

OpenStudy (anonymous):

i did m(45)+(0)(450)=(450)(12), what am i doing wrong ? @Michele_Laino

OpenStudy (michele_laino):

the right equation, is: \[\Large m \times 45 + 0 \times 450 = \left( {m + 450} \right) \times 12\]

OpenStudy (anonymous):

i havve to find 1 m so why is their two

OpenStudy (anonymous):

?

OpenStudy (michele_laino):

since after the impact, te arrow and the apple are ambedded each other, so their momentum is: \[\Large \left( {m + 450} \right) \times 12\;\frac{{{\text{grams}} \times {\text{m}}}}{{{\text{sec}}}}\]

OpenStudy (michele_laino):

after a simple computation, we get: \[\Large m = \frac{{450 \times 12}}{{33}} = ...{\text{grams}}\]

OpenStudy (anonymous):

what happned to the other "m"?

OpenStudy (anonymous):

when we simplify it ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the mass of apple is 450 grams

OpenStudy (anonymous):

i mean the m on the left side

OpenStudy (anonymous):

when we solve for one variable you can't have 2

OpenStudy (michele_laino):

here are more steps: \[\Large \begin{gathered} m \times 45 + 0 \times 450 = \left( {m + 450} \right) \times 12 \hfill \\ 45m = 12m + 450 \times 12 \hfill \\ 33m = 450 \times 12 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ohhh